# Artificial Intelligence for Microbiology and Microbiome Research

### v1

@@ 3.7.3 Disease diagnosis, classification, and clinical outcome prediction
AI can assist in examining novel and intricate data that clinical environments have not fully utilized for diagnostic aims. For instance, for certain diseases involving infections, microbes can generate some VOCs in clinical samples. Hence, we can utilize machine learning to evaluate the odors of those clinical samples to diagnose urinary tract infections [295], active tuberculosis [296], pneumonia [297], and acute exacerbation of chronic obstructive pulmonary disease [298]. For many other diseases associated with disrupted microbiomes, VOCs in clinical samples might not be helpful for disease diagnosis. In this case, we can leverage the microbiome data itself. Indeed, numerous studies have shown microbiome dysbiosis is associated with human diseases [299, 300]. Those diseases include GI disorders, i.e., Clostridioides difficile infection [301], inflammatory bowel disease [302], and irritable bowel syndrome [303], and other non-GI disorders, for example, autism [304], obesity [305], multiple sclerosis [306], hepatic encephalopathy [307], and Parkinson’s disease [308]. Applying supervised classification analysis to the human microbiome data can help us build classifiers that can accurately classify individuals’ disease status, which could assist physicians in designing treatment plans [18].
@@
Phylogenetic tree-based deep learning methods. Classical ML classifiers just treat microbiome data (more specifically, the taxonomic profiles) as regular tabular data, represented as a matrix with rows representing different samples or subjects and columns representing features (i.e., microbial species’ relative abundances). In fact, unlike many other omics, microbial features are endowed with a hierarchical structure provided by the phylogenetic tree defining the evolutionary relationships between those microorganisms. We can exploit the phylogenetic structure and leverage the CNN architecture to deal with species abundance data. With this very simple idea, several deep learning methods (e.g., Ph-CNN [311], PopPhy-CNN [312], taxoNN [313], and MDeep [314]) have been developed. Each method exploits the phylogenetic tree in a slightly different way.
@@
Ph-CNN takes the taxa abundances table and the taxa distance matrix as the input, and outputs the class of each sample [311]. Here, the distance between two taxa is defined as their patristic distance, i.e., the sum of the lengths of all branches connecting the two taxa on the phylogenetic tree. The patristic distance is used together with multi-dimensional scaling to embed the phylogenetic tree in an Euclidean space. Each taxon is represented as a point in Euclidean space preserving the tree distance as well as possible. Since the data is endowed with an intrinsic concept of neighborhood in the input space, Ph-CNN can then use CNN to perform classification. PopPhy-CNN represents the phylogenetic tree and species abundances in a matrix format, and then directly applies CNN to perform classification [312]. taxoNN incorporates a stratified approach to group OTUs into phylum clusters and then applies CNNs to train within each cluster individually [313]. Further, through an ensemble learning approach, features obtained from each cluster were concatenated to improve prediction accuracy. Note that with each phylum cluster, the authors proposed two ways (either based on distance to the cluster center or based on taxa correlations) to order and place correlated taxa together to generate matrix or image-like inputs amenable for CNN. MDeep directly incorporates the taxonomic levels of the phylogenetic tree into the CNN architecture [314]. OTUs on the species level are clustered based on the evolutionary model. This clustering step makes convolutional operation capture OTUs highly correlated in the phylogenetic tree. The number of hidden nodes decreases as the convolutional layer moves forward, reflecting the taxonomic grouping.
@@
Recently, a transformer-based microbial “language” model (MLM) was developed [316]. This MLM was trained in a self-supervised fashion to capture the interactions among different microbial species and the common compositional patterns in microbial communities. The trained MLM can generate robust, context-sensitive representations of microbiome samples to enhance predictive modeling. Note that in this transformer-based MLM, taxa present in each microbiome sample were ranked in decreasing order of abundance to create an ordered list of taxa so that the inputs are analogous to texts. The transformer model then processes these inputs through multiple encoder layers, producing a hidden representation for each taxon. The output of the model includes both sample-level embeddings for classification tasks and context-sensitive embeddings for individual taxon, enabling a nuanced understanding of microbial interactions. By pre-training the transformer using self-supervised learning on large, unlabeled datasets and fine-tuning on specific labeled tasks, this approach leads to improved performance for multiple prediction tasks including predicting IBD and diet patterns.
Note that those three methods (DeepMicro, GDmicro, and the transformer-based MLM) can be applied to any omics data for classification purposes. Their design principles were not based on any unique features of microbiome data.
Despite the development of various methods, a systematical comparison of those deep learning methods and classical machine learning methods on benchmarking datasets is lacking. Since some of those deep learning methods incorporate domain knowledge (i.e., information on the phylogenetic tree, or unlabeled samples), it would be necessary to do that for classical ML methods too, for a fair comparison.
@@
Quite often, we have multi-omics data and clinical data. It would be more insightful to integrate those different data types for better disease status classification or prediction. A straight approach would be to concatenate all datasets into a single view, which is then used as the input to a supervised learning model of choice. A more advanced approach is MOGONET, which jointly explores omics-specific learning using GCNs and cross-omics correlation learning for effective multi-omics data classification [321].
@@ 3.8 Prevention & Therapeutics
3.8.1 Peptides identification & generation
Bacterial resistance to antibiotics is a growing concern. Antimicrobial peptides (AMPs), natural components of innate immunity, are popular targets for developing new drugs. We can divide the AMP activities into different categories, e.g., antibacterial, antiviral, antifungal, antiparasitic, anti-tumor peptides, etc. [325]. Deep learning methods are now commonly adopted by wet-laboratory researchers to screen for promising AMPs. The first work that used neural networks to identify AMPs dates back to 2007, where Lata et al. used a very simple MLP with only one hidden layer [326]. In this work, the authors predicted AMPs based on their N-terminal residues or C-terminal residues, because it has been observed that certain types of residues are preferred at the N-terminal (or C-terminal) regions of the AMPs. In another work published in 2010, Torrent et al. still used a simple MLP with one hidden layer to identify AMPs [327]. In this work, they used the physicochemical properties of AMPs as their features. In total, the authors chosen eight features, including isoelectric point (pI), peptide length, a-helix, b-sheet and turn structure propensity, in vivo and in vitro aggregation propensity and hydrophobicity.
Those early works apparently require quite a lot of domain knowledge and manual feature selection. This effort can be avoided or mitigated by using deep learning models that can automatically learn complex representations and features from raw data, reducing the need for manual feature engineering. For example, in 2018 Veltri et al. proposed a deep neural network model with convolutional and recurrent layers that leverage primary sequence composition [328]. Apparently, it is a hybrid deep learning model. By combining CNN and RNN, the model can extract more meaningful and robust features, avoid the burden of a priori feature construction, and consequently reduce our reliance on domain experts. In 2022, Tang et al. proposed a similar hybrid deep learning model that integrated CNN and RNN [329]. This model is called MLBP: multi-label deep learning approach for determining the multi-functionalities of bioactive peptides. It can predict multi-function, e.g., anti-cancer peptides, anti-diabetic peptides, anti-hypertensive peptides, anti-inflammatory peptides, and anti-microbial peptides, simultaneously. Firstly, the amino acids were converted into natural numbers, and the sequences of all peptides were set to be fixed by using the zero-filled method. Then, an embedding layer was used to learn the embedding matrix of the representation of peptide sequences. The embedding matrix was fed into a CNN to extract the features from the peptide. Then, an RNN is used to analyze streams of the sequence by means of hidden units. Finally, a fully connected layer is applied to the final classification.
The hybrid deep learning approach has been extended further in Ref [330]. The authors started by collecting sequences to build training and test sets and then built and optimized deep learning models to form the AMP prediction pipeline. In particular, the authors included five deep learning models for testing and building the prediction pipeline, including (1) Two CNN + LSTM models; (2) Two CNN + Attention models; and (3) One BERT model. Because the prediction biases were independent of each other, the authors eventually tested the intersection of predictions from various combinations of models (2–5 models). This is a very robust approach. Then they mined metagenomic and metaproteomic data of the human gut microbiome for potential AMPs, further filtering using correlation network analysis between candidate AMPs and bacteria. Finally, they selected promising candidates AMPs from initial screening and further subjected them to efficacy tests against multi-drug resistant (MDR) bacteria, and then in vivo experiments in an animal model. This is a very comprehensive work, clearly demonstrating the power of deep learning models in the identification of AMPs from microbiome data.
Besides identifying natural AMPs, deep learning approaches have also been developed to generate synthetic AMPs. These approaches include GAN and VAE, as well as their conditional variants cGAN and cVAE. The conditional variants enable the generation of peptides satisfying a given condition. For example, AMPGANv2 is based on a bidirectional conditional GAN [331]. It uses generator-discriminator dynamics to learn data-driven priors and control generation using conditioning variables [331]. The bidirectional component, implemented using a learned encoder to map data samples into the latent space of the generator, aids iterative manipulation of candidate peptides. These elements allow AMPGANv2 to generate candidates that are novel, diverse, and tailored for specific applications. Training of GANs was reported to face substantial technical obstacles, such as training instabilities and mode collapse. Hence, VAE-based AMP generations could be an alternative solution. For example, Peptide VAE is based on a VAE, where both encoder and decoder are single-layer LSTMs [332]. The authors also proposed Conditional Latent (attribute) Space Sampling (CLaSS) for controlled sequence generation, aimed at controlling a set of binary (yes/no) attributes of interest, such as antimicrobial function and/or toxicity. HydrAMP is based on a conditional VAE to generate novel peptide sequences satisfying given antimicrobial activity conditions [333]. This method is suitable not only for the generation of AMPs de novo, but also for the generation starting off from a prototype sequence (either known AMPs or non-AMPs).
@@ 3.8.3 Antibiotic discovery
The original MPNN operates on undirected graphs. It is trivial to extend MPNN to directed multigraphs. This yields Directed MPNN, which translates the graph representation of a molecule into a continuous vector via a directed bond-based message passing approach [337]. This builds a molecular representation by iteratively aggregating the features of individual atoms and bonds. The model operates by passing “messages” along bonds that encode information about neighboring atoms and bonds. By applying this message passing operation multiple times, the model constructs higher-level bond messages that contain information about larger chemical substructures. The highest-level bond messages are then combined into a single continuous vector representing the entire molecule.
Stokes et al. discovered a drug halicin by drug repurposing using deep neural networks Chemprop [338, 339] to predict molecules with antibacterial activity. Halicin can against a wide phylogenetic spectrum of pathogens, including Mycobacterium tuberculosis, carbapenem-resistant Enterobacteriaceae, and Clostridioides difficile and pan-resistant Acinetobacter baumannii infections in Murine models [340]. The first module of Chemprop is a local feature encoding function. A molecule’s molecular SMILES string (simplified molecular-input line-entry system) is used as input and transformed into a molecular graph with nodes representing atoms and edges representing bonds using RDKit [341]. The molecular embedding was learned by GCN and was fed into a feed-forward neural networks for classification.
@@
Jame Collins’ lab at MIT recently published two papers on antibiotic discovery [340, 342]. In both papers, they utilized a Direc-MPNN. In principle, their results can be further improved by incorporating a new variant of MPNN, i.e., atom-bond transformer-based MPNN (or ABT-MPNN), which combines the self-attention mechanism in Transformer with MPNNs for better molecular representation and better molecular property predictions. By designing corresponding attention mechanisms in the message-passing and readout phases of the MPNN, ABT-MPNN provides a novel architecture that integrates molecular representations at the bond, atom and molecule levels in an end-to-end way. This model also has a visualization modality of attention at the atomic level, which could be an insightful way to investigate molecular atoms or functional groups associated with desired biological properties, and hence serve as a valuable way to investigate the mechanism of action of drugs (including, but limited to antibiotics).

### v2

@@ 3.7.3 Disease diagnosis, classification, and clinical outcome prediction
AI can assist in examining novel and intricate data that clinical environments have not fully utilized for diagnostic aims. For instance, for certain diseases involving infections, microbes can generate some VOCs in clinical samples. Hence, we can utilize machine learning to evaluate the odors of those clinical samples to diagnose urinary tract infections [401], active tuberculosis [402], pneumonia [403], and acute exacerbation of chronic obstructive pulmonary disease [404]. For many other diseases associated with disrupted microbiomes, VOCs in clinical samples might not be helpful for disease diagnosis. In this case, we can leverage the microbiome data itself. Indeed, numerous studies have shown microbiome dysbiosis is associated with human diseases [405, 406]. Those diseases include GI disorders, i.e., Clostridioides difficile infection [407], inflammatory bowel disease [408], and irritable bowel syndrome [409], and other non-GI disorders, for example, autism [410], obesity [411], multiple sclerosis [412], hepatic encephalopathy [413], and Parkinson’s disease [414]. Applying supervised classification analysis to the human microbiome data can help us build classifiers that can accurately classify individuals’ disease status, which could assist physicians in designing treatment plans [19, 415].
@@
Phylogenetic tree-based deep learning methods. Classical ML classifiers just treat microbiome data (more specifically, the taxonomic profiles) as regular tabular data, represented as a matrix with rows representing different samples or subjects and columns representing features (i.e., microbial species’ relative abundances). In fact, unlike many other omics, microbial features are endowed with a hierarchical structure provided by the phylogenetic tree defining the evolutionary relationships between those microorganisms. We can exploit the phylogenetic structure and leverage the CNN architecture to deal with species abundance data. With this very simple idea, several deep learning methods (e.g., Ph-CNN [190], PopPhy-CNN [191], taxoNN [192], MDeep [193] and PM-CNN [194]) have been developed. Each method exploits the phylogenetic tree in a slightly different way.
@@
Ph-CNN takes the taxa abundances table and the taxa distance matrix as the input, and outputs the class of each sample [190]. Here, the distance between two taxa is defined as their patristic distance, i.e., the sum of the lengths of all branches connecting the two taxa on the phylogenetic tree. The patristic distance is used together with multi-dimensional scaling to embed the phylogenetic tree in an Euclidean space. Each taxon is represented as a point in Euclidean space preserving the tree distance as well as possible. Since the data is endowed with an intrinsic concept of neighborhood in the input space, Ph-CNN can then use CNN to perform classification. PopPhy-CNN represents the phylogenetic tree and species abundances in a matrix format, and then directly applies CNN to perform classification [191]. taxoNN incorporates a stratified approach to group OTUs into phylum clusters and then applies CNNs to train within each cluster individually [192]. Further, through an ensemble learning approach, features obtained from each cluster were concatenated to improve prediction accuracy. Note that with each phylum cluster, the authors proposed two ways (either based on distance to the cluster center or based on taxa correlations) to order and place correlated taxa together to generate matrix or image-like inputs amenable for CNN. MDeep directly incorporates the taxonomic levels of the phylogenetic tree into the CNN architecture [193]. OTUs on the species level are clustered based on the evolutionary model. This clustering step makes convolutional operation capture OTUs highly correlated in the phylogenetic tree. The number of hidden nodes decreases as the convolutional layer moves forward, reflecting the taxonomic grouping. PM-CNN (Phylogenetic Multi-path Convolutional Neural Network) is a phylogeny-based neural network model designed for multi-status classification and disease detection using microbiome data [194]. It organizes microbes according to their phylogenetic relationships and extracts features through a multi-path CNN, a deep learning architecture that processes data via parallel paths. Unlike traditional CNNs, which rely on a single path to extract features layer by layer, potentially leading to information loss due to limited feature representation in complex tasks, the multi-path CNN processes input data through multiple independent paths. This enables feature extraction from diverse angles and scales, enhancing the model’s expressiveness and robustness for complex tasks. In PM-CNN, the the microbial richness features are embedded with phylogeny by multi-layer hierarchical clustering, which are then fed into a multi-path CNN for disease classification. It has been shown that PM-CNN outperforms existing machine learning models in microbiome-based disease classification.
@@
Recently, a transformer-based microbial “language” model (MLM) was developed [196]. This MLM was trained in a self-supervised fashion to capture the interactions among different microbial species and the common compositional patterns in microbial communities. The trained MLM can generate robust, context-sensitive representations of microbiome samples to enhance predictive modeling. Note that in this transformer-based MLM, taxa present in each microbiome sample were ranked in decreasing order of abundance to create an ordered list of taxa so that the inputs are analogous to texts. The transformer model then processes these inputs through multiple encoder layers, producing a hidden representation for each taxon. The output of the model includes both sample-level embeddings for classification tasks and context-sensitive embeddings for individual taxon, enabling a nuanced understanding of microbial interactions. By pre-training the transformer using self-supervised learning on large, unlabeled datasets and fine-tuning on specific labeled tasks, this approach leads to improved performance for multiple prediction tasks including predicting IBD and diet patterns. Multistage Fusion Tabular (MSFT) Transformer [197] is another transformer-based model designed for disease prediction by effectively integrating diverse, high-dimensional tabular data derived from metagenomic data. Its multistage fusion strategy comprises three key modules: an early-stage fusion-aware feature extraction module that enhances the quality of extracted input information, a mid-stage alignment-enhanced fusion module that ensures the retention of critical information during cross-modal learning, and a late-stage integrated feature decision layer that consolidates cross-modal information for robust predictions.
Despite the development of various methods, a systematical comparison of those deep learning methods and classical machine learning methods on benchmarking datasets is lacking. Since some of those deep learning methods incorporate domain knowledge (i.e., information on the phylogenetic tree, or unlabeled samples), it would be necessary to do that for classical ML methods too, for a fair comparison. In addition, while recent deep learning models often outperform traditional machine learning approaches, these advantages typically depend on the availability of large, high-quality labeled datasets and substantial computational resources. In contrast, traditional methods (e.g., random forest, support vector machines) can offer more robust performance in small-sample or low-resource scenarios, where deep learning models are prone to overfitting and require costly training. Thus, model selection should consider the specific data characteristics: deep learning excels with abundant data and computational power, whereas classical machine learning approaches remain valuable for smaller or noisier datasets.
@@
Quite often, we have multi-omics data and clinical data. It would be more insightful to integrate those different data types for better disease status classification or prediction [421]. A straight approach would be to concatenate all datasets into a single view, which is then used as the input to a supervised learning model of choice. A more advanced approach is MOGONET, which jointly explores omics-specific learning using GCNs and cross-omics correlation learning for effective multi-omics data classification [422].
@@ 3.8 Prevention & Therapeutics
3.8.1 Vaccine design
@@ 3.8.3 Antibiotic discovery
The original MPNN operates on undirected graphs. It is trivial to extend MPNN to directed multigraphs. This yields Directed MPNN (or D-MPNN), which translates the graph representation of a molecule into a continuous vector via a directed bond-based message passing approach [435]. This builds a molecular representation by iteratively aggregating the features of individual atoms and bonds. The model operates by passing “messages” along bonds that encode information about neighboring atoms and bonds. By applying this message passing operation multiple times, the model constructs higher-level bond messages that contain information about larger chemical substructures. The highest-level bond messages are then combined into a single continuous vector representing the entire molecule.
D-MPNN has recently been utilized to discover novel antibiotics [436, 437]. For each compound, RDKit generates a graph-based molecular representation from its Simplified Molecular-Input Line-Entry System (SMILES) string. Feature vectors for atoms and bonds are created using computable properties: atom features include atomic number, number of bonds, formal charge, chirality, number of bonded hydrogen atoms, hybridization, aromaticity, and atomic mass; bond features encompass bond type (single, double, triple, or aromatic), conjugation, ring membership, and stereochemistry. During message passing, each bond-associated message (a real number) is updated by summing messages from neighboring bonds, concatenating the current bond’s message with this sum, and applying a neural network layer with a nonlinear activation function. After a fixed number of message-passing steps, the molecule’s messages are aggregated to produce a final molecular representation. This representation is processed through a feed-forward neural network to predict the compound’s activity, such as antibiotic efficacy, cytotoxicity, or proton motive force-altering potential.
@@
The Atom-Bond Transformer-based Message-Passing Neural Network (ABT-MPNN) [438], a novel variant of MPNN, also holds great potential for antibiotic discovery. ABT-MPNN integrates the self-attention mechanism of Transformers with the MPNN framework to enhance molecular representation and improve molecular property predictions. By incorporating tailored attention mechanisms during the message-passing and readout phases, ABT-MPNN provides an innovative end-to-end architecture that unifies molecular representations at the bond, atom, and molecule levels. Additionally, the model offers a visualization modality for attention at the atomic level, providing valuable insights into the molecular atoms or functional groups associated with desired biological properties. This feature makes ABT-MPNN a powerful tool for investigating the mechanisms of action of drugs, including but not limited to antibiotics.
While deep learning approaches effectively identify antibacterial compounds from existing libraries, their structural novelty remains limited. Recently, a deep learning-based framework was developed to design novel antibiotic compounds using either fragment-based or de novo approaches [439]. In the fragment-based approach, D-MPNN screened over 45 million chemical fragments in silico, identifying those with predicted selective antibacterial activity against Neisseria gonorrhoeae and Staphylococcus aureus. These fragments were then expanded into complete molecules using two generative algorithms: a genetic algorithm based on Chemically Reasonable Mutations (CReM) and a VAE. CReM generates new molecules by adding, replacing, or deleting atoms and functional groups starting from a compound of interest. In the de novo approach, no fragment input is required; instead, CReM and VAE design molecules based on patterns learned during training. This framework enables generative deep learning-guided design of antibiotic candidates, facilitating the discovery of novel antibacterial compounds and efficient exploration of vast, uncharted chemical space.
Experimental determination of antimicrobial activity for novel compounds remains time-consuming and costly. While compound-centric deep learning models can accelerate candidate prioritization, they typically require large, custom-labeled training datasets. Recently, a lightweight, data-efficient strategy for antimicrobial discovery was introduced [440]. This approach builds on MolE (Molecular representation through redundancy-reduced Embedding), a self-supervised framework that learns task-agnostic molecular embeddings from unlabeled chemical structures. By leveraging Graph Isomorphism Networks (GINs) and adapting the non-contrastive Barlow-Twins objective to the molecular domain, MolE produces robust, generalizable representations. When integrated with experimentally validated compound–bacteria activity data, these embeddings enable a high-performing XGBoost-based model to predict the antimicrobial potential of any compound. This scalable, cost-effective pipeline significantly accelerates antibiotic discovery.
3.8.4 Antimicrobial peptides identification & design
Bacterial resistance to antibiotics is a growing concern. Antimicrobial peptides (AMPs) are a class of small peptides (containing fewer than 100 amino acids) that widely exist in nature [441]. AMPs are an important part of the innate immune system of different organisms and hold the potential to tackle multidrug-resistant pathogens. In recent years, AMPs have been considered as a substitute for traditiona antibiotics due to their unique antibacterial mechanism.
[…]
Note that those methods work on peptide sequences and are not directly applicable to microbial genomes or metagenomes. Macrel — (Meta)genomic AMP classification and retrieval is the first end-to-end pipeline for the prospection of high-quality AMP candidates from (meta)genomes [449]. In particular, Macrel accepts as inputs paired-end or single-end reads in FastQ format, performs quality-based trimming (using NGLess [450]), assembles contigs (using MEGAHIT [451]), and predicts smORFs (using a modified version of Prodigal [267]). The resulting smORFs (10 100 amino acids) will be used for AMP classification. Marcel combines 6 local and 16 global features, and uses Random Forest for AMP identification. It has been demonstrated that Marcel is comparable to many existing methods with different trade-offs. In particular, Macrel achieves the highest precision and specificity at the cost of lower sensitivity. This property of Macrel (i.e., emphasizing on precision over sensitivity) has been leveraged recently to identify AMPs in the global microbiome (a vast dataset of 63,410 metagenomes and 87,920 prokaryotic genomes from environmental and host-associated habitats) and create the so-called AMPSphere [452], a comprehensive catalog comprising 863,498 nonredundant peptides.
Those AMP prediction methods described above require quite a lot of domain knowledge and manual feature selection. This effort can be avoided or mitigated by using deep learning models that can automatically learn complex representations and features from raw data, reducing the need for manual feature engineering. For example, AMPScanner [453] is a deep neural network model with convolutional and recurrent layers that leverage primary sequence composition. By combining CNN and RNN, AMPScanner can extract more meaningful and robust features and avoid the burden of a priori feature construction. MLBP [205] is another similar hybrid deep learning model that integrated CNN and RNN to predict AMPs. Firstly, the amino acids were converted into natural numbers, and the sequences of all peptides were set to be fixed by using the zero-filled method. Then, an embedding layer was used to learn the embedding matrix of the representation of peptide sequences. The embedding matrix was fed into a CNN to extract the features from the peptide. Then, an RNN is used to analyze streams of the sequence by means of hidden units. Finally, a fully connected layer is applied to the final classification. AMPlify [454] is an attention-enhanced deep learning model that integrates two complementary attention mechanisms (multi-head scaled dot-product attention (MHSDPA) and context attention (CA)) atop a bidirectional LSTM backbone. The bidirectional LSTM first encodes the input sequence recurrently, capturing rich positional and contextual dependencies. The MHSDPA layer then refines this representation by computing multiple weighted views of the sequence using distinct attention heads. Finally, the CA layer produces a single, informative summary vector through a learned weighted average, distilling key contextual insights from the preceding layer. Note that AMPScanner, MLBP, and AMPlify take the peptide sequence as the input. Recently, an end-to-end pipeline for AMP prediction from the human gut microbiome data using deep learning was developed [455]. Instead of using any manually selected features, the authors combined five deep learning models, including (1) Two CNN + LSTM models; (2) Two CNN + Attention models; and (3) One BERT model to automatically extract features for AMP prediction. Because the prediction biases were independent of each other, the authors eventually tested the intersection of predictions from various combinations of those models. This can be considered as an ensemble approach. Their pipeline starts with the prediction of smORFs from 4,409 high-quality representative genomes of microbial organisms present in human microbiome samples. Those genomes were assembled in a previous study [456]. Then they predicted 20,426,401 putative AMPs from the non-redundant smORFs using their ensemble approach. To ensure the predicted AMPs are indeed expressed, they leveraged multiple metaproteome datasets to further identify 2,349 candidate AMPs. Finally, they performed correlation analysis between the abundances of those candidate AMPs and bacteria, using metagenomic datasets from 15 independent large cohorts. They hypothesized that those candidate AMPs with strong negative correlations with members of a microbiome are more likely to be functional. Although this study focused on the human gut microbiome, the developed pipeline can be easily adopted for AMP mining in other microbiomes.
Molecular de-extinction (MDE) is the resurrection of extinct biomolecules that are no longer encoded by living organisms. MDE is driven by the hypothesis that molecules conferring evolutionary advantages to extinct species may offer benefits in the modern global environment. Recently, MDE has emerged as a novel framework for AMP discovery by expanding the therapeutic search space through paleoproteome mining. Specifically, panCleave [457], a Random Forest-based model, was developed to perform proteome-wide cleavage site prediction and computational proteolysis (i.e., the in silico digestion of human proteins into peptide fragments) to prospect AMPs encrypted within extinct and extant human proteins. In vitro antimicrobial activity was confirmed for both modern and archaic protein-derived peptides identified by panCleave. Lead candidates demonstrated proteolytic resistance and variable membrane permeabilization. Recently, a deep learning method APEX has been developed to mine proteomes of extinct species [458] and archaeal organisms [459] for the discovery of AMPs. APEX utilized a hybrid of recurrent and attention neural networks to extract peptide sequence information, which was further processed by two fully connected neural networks to predict species-specific antimicrobial activities (i.e., a multi-output regression task) or general antimicrobial or not (i.e., a binary classification task), respectively. To improve the prediction performance, an ensemble learning approach was adopted by selecting the top eight APEX models (with different neural network architectures and training strategies). The guidelines of using machine learning to mine genomes and proteomes for AMP discovery have been summarized nicely in a tutorial [460].
Besides identifying natural AMPs, deep learning approaches have also been developed to design synthetic AMPs. Early attempts leveraged GAN and VAE, as well as their conditional variants cGAN and cVAE. For example, AMPGANv2 is based on a bidirectional conditional GAN [206]. It uses generator-discriminator dynamics to learn data-driven priors and control generation using conditioning variables [206]. The bidirectional component, implemented using a learned encoder to map data samples into the latent space of the generator, aids iterative manipulation of candidate peptides. These elements allow AMPGANv2 to generate candidates that are novel, diverse, and tailored for specific applications. Training of GANs was reported to face substantial technical obstacles, such as training instabilities and mode collapse. VAE-based AMP generations could be an alternative solution. For example, Peptide VAE is based on a VAE, where both encoder and decoder are single-layer LSTMs [461]. The authors also proposed Conditional Latent (attribute) Space Sampling (CLaSS) for controlled sequence generation, aimed at controlling a set of binary (yes/no) attributes of interest, such as antimicrobial function and/or toxicity. HydrAMP is based on a conditional VAE to generate novel peptide sequences satisfying given antimicrobial activity conditions [462]. This method is suitable not only for the generation of AMPs de novo, but also for the generation starting off from a prototype sequence (either known AMPs or non-AMPs). Recently, a conditional denoising VAE (cdVAE) model was developed for AMP generation [463]. For each AMP sample, 10 physicochemical properties—molecular weight, isoelectric point, GRAVY, aromaticity, instability index, disulfide bonds, molecular volume, and secondary structure fractions ($\alpha$-helix, $\beta$-sheet, and random coil)—are computed and normalized to form a property vector. The AMP sequence is one-hot encoded and passed through an embedding layer to generate an embedding vector. Gaussian noise from a normal distribution is added to this embedding vector to create noised samples. These noised samples, along with the property vector, are processed through the cdVAE’s encoder and decoder, producing a reconstructed sample that achieves a denoising effect. Specifically, the encoder maps the AMP input data distribution to a standard normal distribution, while the decoder transforms samples from this distribution into AMP sequences with desired physicochemical properties. Notably, the property vector guides the VAE training process, enabling the decoder to map the standard normal distribution to a conditional distribution of AMPs with tailored physicochemical characteristics.
EvoGradient [207] (Evolution guided by Gradient) is an interesting attempt to virtually modify peptide sequences to produce more potent AMPs, akin to in silicon directed evolution. This was achieved through an iterative process that combines gradient descent with projection to find a locally optimal sequence near the original peptide. In each step, a scaled gradient is subtracted from the input vector, followed by a projection back to one-hot format to check for any amino acid mutations. If a mutation is found, the projected result is used for the next iteration; otherwise, projection is skipped. This amino acid editing is performed iteratively until half of the original peptide’s amino acids are altered or a local optimum is reached. deepAMP [208], which takes peptides with low antimicrobial activity as inputs and output analogs with high antimicrobial activity and broad-spectrum resistance. It achieves this through a pre-training and multi-stage fine-tuning strategy, coupled with a sequence degradation approach for data augmentation. Promising results demonstrate deepAMP’s ability to accelerate the discovery of effective AMPs against drug-resistant bacteria, including Staphylococcus aureus, Klebsiella pneumoniae, and Pseudomonas aeruginosa. AMPainter [464] is a deep reinforcement learning framework that unifies optimization and generation of AMPs in a single pipeline. It is applied to three peptide classes: known AMPs, signal peptides (SPs), and random sequences. AMPainter outperforms ten benchmark models in enhancing predicted antimicrobial potency and sequence diversity when optimizing known AMPs. Given a set of input peptide sequences, AMPainter iteratively evolves them into improved AMPs through a three-step process: (1) a policy network identifies mutation sites, (2) a fine-tuned language model proposes residue replacements, and (3) HyperAMP—a hypergraph neural network pre-trained on known AMPs and their activity labels—evaluates antimicrobial potential. The predicted activity scores serve as rewards to update the policy network via reinforcement learning, enabling continuous improvement across iterations.
TransSAFP [465] is a deep learning model that predicts the functional activity of self-assembling functional peptides (SAFPs) with minimal experimental annotation. This enables de novo design of AMPs. TransSAFP comprises two modules: a pretraining module and a transfer learning module. The pretraining architecture is a multi-head self-attention transformer with a cross-attention layer, while the downstream transfer learning module appends a self-attention block to the pretrained latent representation. The resulting model designed a lead self-assembling peptide that demonstrates excellent in vivo therapeutic efficacy against intestinal bacterial infections in mice, with superior biofilm eradication and no induction of acquired drug resistance. Mechanistic studies reveal that the peptide self-assembles on bacterial membranes to form nanofibrous structures, effectively killing multidrug-resistant pathogens. TransSAFP is promising for the customized design of functional peptide materials.